# Learning nuclear cross sections across the chart of nuclides with graph neural networks

### v2

@@ III.2 Representation Learning
Representation learning aims to condense raw input data into a concise format, preserving crucial information for target tasks. Ideally, the learned representation provides a lower-dimensional space which makes subsequent learning tasks more efficient. In our setup, the goal of representation learning is to learn the space of all cross sections. By doing so, not only can all known cross sections of a nucleus be encoded in the representation space, we can also predict new cross sections for known and unknown nuclei. This prediction task then just translates into the task of finding the right vector in the learned representation space.
Figure 2: (a) Variational Autoencoder framework illustrating the encoding and decoding process with a latent space bottleneck. (b) Implicit Neural Representation (INR) architecture integrated with a hypernetwork (HyperNet). The hypernetwork generates weights for the INR, informed by a convolutional encoder that processes input data. Note that the INR takes input coordinates (representing the energy index with a length of 256) and maps them to the corresponding cross-sections.
@@ III.2.1 Variational Autoencoder
One of the most commonly used approaches for representation learning is an autoencoder (AE) architecture. An AE is a neural network designed to encode input data into a lower-dimensional space (i.e., bottleneck) and then decode it back to the original data space. As illustrated in Fig. 2(a), the autoencoding objective is centered around minimizing the discrepancy between the original input and the reconstructed output. For variational autoencoders (VAEs) [25], the standard autoencoding objective is augmented by the inclusion of variational inference [33], where the encoding process maps input data not to a fixed point but to a probability distribution in the latent space. This dual training objective involves both minimizing the reconstruction error and aligning the latent space distribution with a specified form, i.e., a standard Gaussian distribution (with parameter $\sigma$ and $\mu$). Specifically, the VAE training loss can be defined by two key terms, the reconstruction loss and the Kullback-Leibler (KL) divergence,
@@
The first term encourages the VAE to accurately reconstruct the original input data. It measures the negative log-likelihood of the decoder distribution ($p_{\theta}(x|z)$) generating data that resembles the input ($x$) given a sampled latent code ($z$) from the encoder distribution ($q_{\phi}(z|x)$). The second term acts as a regularizer, forcing the learned latent distribution ($q_{\phi}(z|x)$) to be similar to a prior distribution ($p(z)$), i.e., a standard normal Gaussian. The hyperparameter $\lambda$ controls the balance between reconstruction accuracy and the regularity of the latent space. Following the original VAE formulation, we set $\lambda$ to $1$.
@@
In this work, the encoder of the VAE model comprises a sequence of convolutional layers followed by ReLU activation functions [34] and a flattening operation. This architecture is designed to progressively reduce the spatial dimensions of the inputs while increasing the number of feature channels. Our encoder includes four convolutional layers – noted conv_1, conv_2, conv_3, and conv_4 in Fig. 2 – with increasing filters (32, 64, 128, 256), kernel size $4$, and stride $2$. Each layer is followed by a ReLU activation. We then convert the output of convolutional layers into a flattened vector of size $k$.
The decoder comprises a sequence of transposed convolutional layers aimed at reconstructing the original inputs from the latent space. It also employs ReLU activations and concludes with a sigmoid activation for the reconstructed inputs. As the mirror opposite of the encoder, the decoder progressively refines the representation through four layers with specific filter sizes and strides: 128 filters with a kernel size of 5 and stride of 2, followed by 64 filters with a kernel size of 5 and stride of 2, then 32 filters with a kernel size of 6 and stride of 2, and finally, the output layer with a kernel size of 24 and a stride of 8. The output layer returns a vector of size $m$. By jointly optimizing the encoder and decoder, the VAE learns a compact and informative representation of given inputs.
@@ III.2.2 Implicit Neural Representation
INR is a method for representing and modeling input data without explicitly defining the geometry or structure of objects [35, 32]. Traditional data encoding methods rely on discrete representations, such as data grids, which are constrained by their spatial resolution and are susceptible to inherent discretization artifacts. In contrast, INR represents input data (including sound, time-series, images, and 3D scenes) as a continuous function of its coordinate locations, where each value is generated independently. This function is approximated by a deep neural network. INR offers flexibility and adaptability for encoding data independent of resolution through the use of coordinate grids, making them highly effective with numerous potential applications [36, 35, 37].
@@
As shown in Fig. 2(b), our framework integrates a hypernetwork with an INR. The convolutional encoder leverages the same configuration as in Fig. 2(a), featuring a series of convolutional layers. It processes cross-section data and compresses it into a 32-dimensional latent vector. This vector is then passed to the hypernetwork decoder, which consists of standard multi-layer perceptron layers with ReLU activations. The hypernetwork’s output provides the weights for the INR. Finally, the INR takes coordinates as input and reconstructs the original input data. This pipeline enables networks to be trained in an end-to-end manner by minimizing the following loss function
[…]
where $\mathcal{L}_{\rm recon}$, $\mathcal{L}_{\rm latent}$, and $\mathcal{L}_{\rm weights}$ represent the reconstruction loss, regularization terms for latent vectors and weights in the hypernetwork. The hyperparameters $\alpha$ and $\beta$ are utilized to balance the contributions of the different losses. Specifically, each loss term is computed as follows:
[…]
@@ III.3 Graph Neural Networks
GNNs are optimizable transforms that operate on all attributes of a graph (nodes, edges, global context, inherent relationships, etc.) while preserving symmetries and the underlying structure of the data. This lends to their excellent capabilities to estimate non-Euclidian sub-spaces and hence, GNNs are ideal for this case where we seek to expose and explore the relationships between nuclear cross-sections across a two-dimensional grid. GNNs accept a graph as input (the encoded 2D TENDL grid) with the information contained in the nodes (feature vectors) and perform transformations without changing the connectivity of the input. GNNs also produce a graph at the output.
The graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is created using the encoded data obtained from the encoding module of the representation learning architecture (VAE or INR) with a mask imposed on the test set so that it is not visible during training. For optimization, we calculate the loss as
$\mathcal{L}_{\rm GNN}=\|\chi(\mathcal{G})-\mathbf{y}_{z}\|_{2}\,,$ (4)
where $\mathcal{G}$ is the input graph to the GNN, $\chi$ is the chosen GNN architecture, and $\mathbf{y}_{z}$ is the encoded cross section representation from the TENDL library, i.e. the encoded ground truth. The loss optimization is performed in the latent space to ensure better convergence characteristics and to leverage the excellent performance of the representation learning module; see Fig. 3.
In this work, GNNs are constructed using the “message passing neural network” (MPNN) backbone [38]. MPNNs follow an iterative scheme to update the nodes by aggregating information from other nearby nodes. The aggregation function is permutation invariant and passes on the aggregated information to the next layer. For the regular grid graph considered in this paper, the forward pass through the GNN has two phases – a message passing phase and a readout phase. In the message passing phase for a particular node, the message function $M_{t}$ aggregates the information from its neighbors and updates the node according to some update function $U_{t}$. At any time $t$, the hidden state $h_{i}^{t}$ at node $v_{i}$ is updated as
@@
where $N(v_{i})$ denotes the neighbors of node $v_{i}$ in the graph $\mathcal{G}$. Here, neighbors are defined by the adjacency to the original node: the neighbors of $v_{i}$ are all the nodes that are next to it in the $(Z,N)$ grid. The readout phase computes a feature vector for the whole graph using some readout function $R_{t}$. $M_{t},U_{t}$, and $R_{t}$ are all learnable differentiable functions that are separately estimated using machine learning methods.
We construct three different GNN architectures based on ARMAConv [39], GatedConv [40], and the traditional GCN (Graph Convolution Network) [41] with the goal of establishing a performance baseline for this data type. The traditional GCN employs a convolution layer that operates on the input feature matrix $\mathbf{X}$, with the adjacency matrix $\mathbf{A}$, a degree matrix $\mathbf{D}$, and filter parameters $\Theta$ in the Fourier domain. The positional relationships (edge $e_{ij}$ between any two input nodes $v_{i},v_{j}$) of the input data are captured by $\mathbf{A}$. In contrast, GatedConv is designed for sequence prediction using the recurrence of the basic graph propagation model where a sequence of node annotations $\mathcal{X}$ and node state machines $\mathcal{H}$ are used to produce an output sequence. Finally, ARMAConv experiments explore the effect of auto-regressive filter layers instead of the traditional polynomial spectral filters. This layer can model complex relationships with fewer parameters and learns a representation independent of the underlying graph architecture. The initial node features $\mathbf{X}$ are transformed using learnable parameters $\mathbf{W}$ and $\mathbf{V}$, the original Laplacian of the graph, $\mathbf{L}$, and some non-linearity $\sigma(\cdot)$ such as ReLU. All these architectures considered use the MPNN backbone with slight variations either in terms of the type of filter used or the node update method.
@@ IV Results
We recall that our proposed framework consists of two main components. In a first step, we learn the latent vector encoding the nuclear cross sections with dimensionality reduction techniques. In a second step, we train a GNN to predict the latent vector in neighboring nuclei by leveraging the learned latent vector before the decoder transforms the data back to the original 256-dimension cross-section. Since a crucial part of dealing with GNNs is the lower-dimension representation space, we first analyze our two encoding mechanisms – the VAE model of Fig. 2(a) and the INR model of Fig. 2(b). We then present the results of several experiments designed to study the behavior of different GNN architectures. Specifically our goal is to infer unknown (i.e., test set) cross-sections based on the knowledge learned from the known cross-sections in the training set.
@@ IV.1 Encoding Performance
The quality of the latent representation plays a critical role to predict accurate cross sections. Therefore, in this section we focus on evaluating the quality of the latent space learned by both the VAE and INR.
@@
Figure 4: Mean prediction errors calculated using Eq. (8) of the GNN for each nucleus in the TENDL dataset. For better visualization, all plots are normalized between $\epsilon_{\rm min}$ and $0.1\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error over all plots. The three panels on the left side are based on VAE encoding while the three on the right on INR encoding. ARMA, Gated and GCN refer to the three types of GNNs considered in this work. The red square corresponds to the windowed test data: the GNN was not trained on it.
For this experiment, we randomly split the dataset containing $2151$ samples into a training set ($\approx$90%) and a testing set ($\approx$10%). To ensure reproducibility and fair comparisons, we employed a fixed random seed, guaranteeing both models were trained on identical datasets. Prior to model training, we normalized the input data by dividing each cross-section by its corresponding maximum value across the entire dataset. This ensures that each cross-section falls within the range of $[0,1]$. Then, the VAE model was trained on a multi-channel cross-section dataset for $5,000$ epochs with a batch size of $32$. We utilized the Adam optimizer (initial learning rate $1\times 10^{-3}$) for optimization. The INR model, trained on the same dataset, also used the Adam optimizer (learning rate $5\times 10^{-5}$) but for $600$ epochs. Notably, the INR model demonstrated significantly faster convergence compared to the VAE training while maintaining high quality of latent representations.
Figure 3 illustrates the typical performance of both the VAE (top) and INR (bottom) for a nucleus within the TENDL dataset. Each color represents one of the four cross-section channels considered in this work. The curves with symbols (square, circle, x, inverted triangle) represent the reconstructed inputs generated by either the VAE or the INR, whereas plain curves depict the ground truth inputs. As seen in this figure, both models effectively reconstruct given test samples, closely aligning with the ground truth. These results demonstrate that both models are able to encode the multi-channel cross-section input into a compact, 32-dimensional latent vector with minimal loss in reconstruction accuracy.
To quantify reconstruction performance, we adopt the Peak Signal-to-Noise Ratio (PSNR). The PSNR is defined as follows:
@@
where $\text{MSE}=\frac{1}{n}\sum^{n}_{i=1}\|\mathbf{y}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2}$, $\hat{\mathbf{x}}_{i}$ represents the reconstructed input generated by either the VAE or the INR, while $\mathbf{y}_{i}$ denotes the original cross-sectional inputs. Here, $n$ is the total number of testing samples. MAX represents the maximum possible value of the signal, which is equal to $1$. We evaluate each model using this metric on the entire test set. The VAE achieves a score of $43.67$, while the INR achieves a higher score of $47.20$. Overall, the performance of the INR is significantly better as this model is better able to precisely capture high-frequency information such as thresholds, for example the fact that the cross section for the $(n,2n)$ cross section is identically 0 unless the energy of the incident neutron exceeds the neutron separation energy. For completeness, Appendix VII.1 shows the evolution of the train and test losses as well of the train and test PSNR as a function of epochs.

### v3

@@ III.2 Representation Learning
Representation learning aims to condense raw input data into a concise format that preserves crucial information. Ideally, the learned representation provides a lower-dimensional space which makes subsequent learning tasks more efficient. In our setup, the goal of representation learning is to learn the latent space of all cross sections. By doing so, not only can all known cross sections of a nucleus be encoded in the latent space, we can also predict new cross sections for known and unknown nuclei. This prediction task then just translates into the task of finding the right vector in the learned representation space.
Figure 2: Panel (a): Variational Autoencoder framework illustrating the encoding and decoding process with a latent space bottleneck. Panel (b): Implicit Neural Representation (INR) architecture integrated with a hypernetwork (HyperNet). The hypernetwork generates weights for the INR, informed by a convolutional encoder that processes input data. Note that the INR takes input coordinates, $\{e_{i}\}$ (representing the energy index with a length of $m=256$) and maps them to the corresponding cross sections.
@@ III.2.1 Variational Autoencoder
One of the most commonly used approaches for representation learning is an autoencoder (AE) architecture. An AE is a neural network designed to encode input data into a lower-dimensional space (i.e., bottleneck) and then decode it back to the original data space. The autoencoding objective is centered around minimizing the discrepancy between the original input and the reconstructed output. For variational autoencoders (VAEs) [25], the standard autoencoding objective is augmented by the inclusion of variational inference [33], where the encoding process maps input data not to a fixed point but to a probability distribution in the latent space. This dual training objective involves both minimizing the reconstruction error and aligning the latent space distribution with a specified form, i.e., a standard Gaussian distribution (with parameters $\mu$ and $\sigma$), as illustrated in Fig. 2(a). Specifically, the VAE training loss can be defined by two key terms: the reconstruction loss and the Kullback-Leibler (KL) divergence,
@@
The first term encourages the VAE to accurately reconstruct the original input data. It measures the negative log-likelihood of the decoder distribution ($p_{\theta}(x|z)$) generating data that resembles the input ($x$) given a sampled latent code ($z$) from the encoder distribution ($q_{\phi}(z|x)$). The second term acts as a regularizer, forcing the learned latent distribution ($q_{\phi}(z|x)$) to be similar to a prior distribution ($p(z)$), i.e., a standard normal Gaussian. The hyperparameter $\lambda$ controls the balance between reconstruction accuracy and the regularity of the latent space. In this work, we set $\lambda$ to $0.1$.
@@
In this work, the encoder of the VAE model comprises a sequence of convolutional layers followed by Rectified Linear Unit (ReLU) activation functions [34] and a flattening operation. This architecture is designed to progressively reduce the spatial dimensions of the inputs while increasing the number of feature channels. Our encoder includes four convolutional layers —noted conv_1, conv_2, conv_3, and conv_4 in Fig. 2(a)— with increasing filters (32, 64, 128, 256), kernel size 4, and stride 2. We then convert the output of convolutional layers into a flattened vector of size $k$.
The decoder comprises a sequence of transposed convolutional layers aimed at reconstructing the original inputs from the latent space. It also employs ReLU activations and concludes with a sigmoid activation for the reconstructed inputs. As the mirror opposite of the encoder, the decoder progressively refines the representation through four layers with specific filter sizes and strides: 128 filters with a kernel size of 5 and stride of 2, followed by 64 filters with a kernel size of 5 and stride of 2, then 32 filters with a kernel size of 6 and stride of 2, and finally, the output layer with a kernel size of 24 and a stride of 8. The output layer returns a vector of size $m$. By jointly optimizing the encoder and decoder, the VAE learns a compact representation of given inputs.
@@ III.2.2 Implicit Neural Representation
INR is a method for representing and modeling input data without explicitly defining the geometry or structure of objects [35, 32]. Traditional data encoding methods rely on discrete representations, such as data grids, which are constrained by their spatial resolution and are susceptible to inherent discretization artifacts. In contrast, INR represents input data (including sound, time-series, images, and 3D scenes) as a continuous function of its coordinate locations, where each value is generated independently. This function is approximated by a deep neural network. INR offers flexibility and adaptability for encoding data independently of resolution through the use of coordinate grids, making them highly effective with numerous potential applications [36, 35, 37].
@@
As shown in Fig. 2(b), our framework integrates a hypernetwork with an INR. The convolutional encoder leverages the same configuration as in Fig. 2(a), featuring a series of convolutional layers. It processes cross section data and compresses it into a 32-dimensional latent vector. This vector is then passed to the hypernetwork decoder, which consists of standard multi-layer perceptron layers with ReLU activations. The hypernetwork’s output provides the weights for the INR. Finally, the INR takes coordinates as input and reconstructs the original input data. This pipeline enables networks to be trained in an end-to-end manner with the following loss function:
$\mathcal{L}_{\rm INR}=\mathcal{L}_{\rm recon}+\alpha\mathcal{L}_{\rm latent}+\beta\mathcal{L}_{\rm weights},$ (2)
where $\mathcal{L}_{\rm recon}$, $\mathcal{L}_{\rm latent}$, and $\mathcal{L}_{\rm weights}$ represent the reconstruction loss, regularization terms for latent vectors, and weights in the hypernetwork, respectively. The hyperparameters $\alpha$ and $\beta$ are utilized to balance the contributions of the different losses. Specifically, each loss term is computed as follows:
$\mathcal{L}_{\rm recon}$ $=\frac{1}{N_{\rm batch}}\sum_{i=1}^{N_{\rm batch}}\|\mathbf{y}_{i}-\hat{\mathbf{x}}_{i}\|_{2},$ (3a)
$\mathcal{L}_{\rm latent}$ $=\frac{1}{N_{\rm batch}}\sum_{i=1}^{N_{\rm batch}}\|\mathbf{z}_{i}\|_{2}^{2},$ (3b)
$\mathcal{L}_{\rm weights}$ $=\frac{1}{N_{\rm batch}}\frac{1}{N_{\mathcal{W}}}\sum_{i=1}^{N_{\rm batch}}\sum_{\mathbf{w}\in\mathcal{W}}\|\mathbf{w}_{i}\|_{2}^{2}\,,$ (3c)
@@ III.3 Graph Neural Networks
GNNs are optimizable transforms that operate on all attributes of a graph (nodes, edges, global context, inherent relationships, etc.) while preserving symmetries and the underlying structure of the data. Hence, GNNs are well suited to expose and explore the relationships between nuclear cross sections across a two-dimensional grid—the nuclear chart. GNNs accept a graph as input (for example, the encoded 2D TENDL grid) with the information contained in the nodes (feature vectors) and perform transformations without changing the connectivity of the input. GNNs also produce a graph as the output.
In this study, we are interested in graph-based imputation, where the goal is to predict missing node features by leveraging the structural and feature information from neighboring nodes. Two primary learning paradigms are commonly considered, transductive and inductive settings. In the transductive setting, the model is trained and evaluated on the same graph, where the node and edge structure remain fixed. In the inductive setting a subset of the graph is used for training while testing is performed on a disjoint subset or entirely new graph. Such a paradigm generally requires much stronger generalization power of the network. In this initial work, we focus on the transductive setting, as our primary interest lies in improving predictions by fully exploiting structural correlations across neighboring nuclei within the given nuclear chart.
III.3.1 Graph Connectivity and Message-Passing
The graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is constructed by representing each nucleus as a node. The node features are the learned latent representation of the nuclear cross sections obtained from the encoding module (either VAE and INR), i.e., vectors of dimension $k=32$. The edges $\mathcal{E}$ define the structural relationship between nuclei based on their positions in the $(Z,N)$ chart. Here, we employ 2-skip edge connectivity, where each node skips its 1-hop and 2-hop neighbors and connects to nodes located three steps away along the horizontal and vertical directions. This design choice is motivated by our goal to impute a relatively large contiguous missing region (e.g., a $9\times 9$ test window). Using only local connectivity would limit the model’s ability to propagate information into the interior of the masked test region. The designed skip connectivity enables longer-range message passing and allows the model to capture broader trends in the nuclear chart. We identified a performance trade-off between local dense connectivity and skip connectivity, particularly with respect to the size of the test window. For example, denser edge connections tend to yield better performance for smaller test windows, whereas a noticeable performance drop is observed when applied to larger test windows (e.g., $9\times 9$), where longer-range dependencies become more critical.
To effectively model such neighbor-dependent interactions, our GNN architecture $\chi(\cdot)$ leverages the message passing neural network (MPNN) framework [38], where the hidden state of each node is iteratively updated by aggregating information from its neighbors. Specifically, for the regular grid graph considered in this paper, the forward pass through the GNN has two phases—a message passing phase and a readout phase. In the message passing phase for a particular node, the message function $M_{t}$ aggregates the information from its neighbors and updates the node according to some update function $U_{t}$. At any time $t$, the hidden state $h_{i}^{t}$ at node $v_{i}$ is updated as
@@ III.3.1 Graph Connectivity and Message-Passing
where $N(v_{i})$ denotes the neighbors of node $v_{i}$ in the graph $\mathcal{G}$, and $e_{ij}$ represents an optional edge attributes.
III.3.2 GNN Training
In our GNN-based imputation framework, training is performed by randomly masking a subset of nodes at each iteration and replacing their features with random placeholders. The GNN is then tasked with predicting the original values in the masked regions using the features of neighboring nodes and the graph structure. Importantly, the loss is computed only over the masked nodes. The test nodes (held-out $9\times 9$ region centered at $(Z_{0},N_{0})=(42,60)$) are also replaced with fixed random placeholders, using a consistent random seed, and are excluded from training loss computations.
We adopt two different training strategies depending on the underlying representation model, VAE or INR, used to generate node features. For VAE-based embeddings, we found that training the GNN to minimize reconstruction loss (Mean Square Error) in the original cross section space, i.e., between decoded predictions and ground truth, works well, particularly when the decoder is fine-tuned jointly with the GNN. This end-to-end training allows the decoder to adapt to the GNN’s latent predictions, resulting in more accurate predictions. By contrast, for INR-based embeddings, we observe that freezing the INR decoder and minimizing the loss directly in the latent space leads to better performance. This might be because the INR decoder with hypernetwork that generates the weights of an implicit neural function makes it difficult to fine-tune effectively. Formally, the two training objectives read
$\mathcal{L}_{\text{GNN/VAE}}$ $=\frac{1}{|\mathcal{M}|}\sum_{v_{i}\in\mathcal{M}}\left\|\phi\left(\chi(\mathcal{G})_{i}\right)-\mathbf{y}_{i}\right\|^{2}\,,$ (6)
$\mathcal{L}_{\text{GNN/INR}}$ $=\frac{1}{|\mathcal{M}|}\sum_{v_{i}\in\mathcal{M}}\left\|\chi(\mathcal{G})_{i}-\mathbf{z}_{i}\right\|^{2}\,.$ (7)
Here, $\mathcal{M}$ denotes the set of masked training nodes. $\chi(\mathcal{G})_{i}$ denotes the GNN-predicted latent vector for node $v_{i}$, and $\phi$ represents the decoder that maps the latent vectors to the original cross section domain. $\mathbf{y}_{i}$ is the ground-truth cross section, while $\mathbf{z}_{i}$ refers to the corresponding ground-truth latent representation. The overall imputation pipeline is summarized in Algorithm 1. Furthermore, we refer the reader to Appendix VII.3 for a detailed comparisons for training strategies under different loss spaces and decoder fine-tuning options.
Algorithm 1 Graph-Based Imputation on Embeddings
1:Node feature extractor $\mathrm{E}$ (VAE or INR encoder), decoder $\phi$ (VAE or INR), GNN $\chi$, training set $\mathcal{T}$, test region $\mathcal{S}$ (e.g., $9\times 9$ window), mask probability $p=10$%, number of epochs $T$
2:Predicted node features for $\mathcal{S}$
3:Encode raw cross section data: $\mathbf{z}_{i}\leftarrow\mathrm{E}(\mathbf{y}_{i})$ for all $v_{i}\in\mathcal{T}\cup\mathcal{S}$
4:Build graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with chosen edge connectivity (e.g., 2-skip edge connectivity)
5:Replace node features of all $v_{j}\in\mathcal{S}$ with fixed random noise: $\mathbf{z}_{j}\leftarrow\eta_{j}$ with fixed seed
6:for $t=1$ to $T$ do
7:     Sample a random subset $\mathcal{M}\subset\mathcal{T}$ with $|\mathcal{M}|=p\cdot|\mathcal{T}|$
8:     Replace node features of $\mathcal{M}$ with random noise $\eta_{i}$
9:     $\hat{\mathbf{z}}_{i}\leftarrow\chi(\mathcal{G})_{i}$ for all $v_{i}\in\mathcal{V}$
10:     if using VAE then
11:         $\hat{\mathbf{y}}_{i}\leftarrow\phi(\hat{\mathbf{z}}_{i})$triangleright Decode latent
12:         $\mathcal{L}\leftarrow\frac{1}{|\mathcal{M}|}\sum_{v_{i}\in\mathcal{M}}\left\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\right\|^{2}$triangleright Eq. (6)
13:     else
14:         $\mathcal{L}\leftarrow\frac{1}{|\mathcal{M}|}\sum_{v_{i}\in\mathcal{M}}\left\|\hat{\mathbf{z}}_{i}-\mathbf{z}_{i}\right\|^{2}$triangleright Eq. (7)
15:     end if
16:     Backpropagate and update parameters of $\chi$ (and $\phi$ if decoder is fine-tuned)
17:end for
18:Use trained $\chi$ (and $\phi$) to predict missing features for all $v_{j}\in\mathcal{S}$
19:Return $\hat{\mathbf{y}}_{j}\leftarrow\phi(\chi(\mathcal{G})_{j})$
III.3.3 GNN Architectures
We evaluate three types of message-passing neural network architectures, ARMAConv [39], GatedConv [40], and GraphSAGE [41]. The ARMA convolutional layer simulates graph convolution via approximated rational filters using a stack of recurrent units. It supports deeper feature propagation and has shown stability in capturing long-range dependencies. With shared weights across layers, it balances model expressiveness and parameter efficiency. The GatedConv integrates recurrent updates with gated mechanisms, allowing information to flow adaptively over multiple time steps. The gates help control feature propagation, which can be useful for graphs where information needs to be selectively passed or retained. The GraphSAGE generates node embeddings by aggregating features from a node’s neighborhood. It supports scalable learning and provides robust performance across a variety of graph learning tasks. In our setup, it achieves the highest imputation performance with VAE embeddings while maintaining parameter efficiency. All these architectures considered use the MPNN backbone with slight variations either in terms of the type of filter used or the node update method, and each architecture is trained under the same data and masking setup.
@@ IV Results
We recall that our proposed framework consists of two main stages. First, we learn low-dimensional latent representations of nuclear cross sections using dimensionality reduction techniques. Second, we train GNN to impute the missing node features (i.e., latent vectors) in the nuclear chart. Since the effectiveness of GNNs in this task depends on the quality of embedding space, we begin by analyzing the two types of encoders used in this study, the VAE model of Fig. 2(a) and the INR model of Fig. 2(b). We then evaluate the performance of several GNN architectures for the downstream imputation task, where the goal is to predict unknown cross sections in the test region using information learned from known regions in the training set.
@@ IV.1 Encoding Performance
To train models for latent representations, we randomly split the dataset containing $2171$ samples into a training set ($\approx$90%) and a testing set ($\approx$10%). To ensure reproducibility and fair comparisons, we employed a fixed random seed, guaranteeing both models were trained on identical datasets. Prior to model training, we normalized the input data by dividing each cross section by its corresponding maximum value across the entire dataset. This ensures that each cross section falls within the range of $[0,1]$. Then, the VAE model was trained on a multi-channel cross section dataset for $5,000$ epochs with a batch size of $32$. We utilized the Adam optimizer (initial learning rate $1\times 10^{-3}$) for optimization. The INR model, trained on the same dataset, also used the Adam optimizer (learning rate $5\times 10^{-5}$) but for $600$ epochs. Notably, the INR model demonstrated significantly faster convergence compared to the VAE training while maintaining high quality of latent representations.
@@
Figure 3 illustrates the performance of both the VAE (top) and INR (bottom) for a nucleus within the TENDL dataset. Each color represents one of the four cross section channels considered in this work. The curves with symbols (square, circle, x, inverted triangle) represent the reconstructed inputs generated by either the VAE or the INR, whereas plain curves depict the ground truth inputs. As seen in this figure, both models effectively reconstruct given test samples, closely aligning with the ground truth. These results demonstrate that both models are able to encode the multi-channel cross section input into a compact, 32-dimensional latent vector with minimal loss in reconstruction accuracy.
Figure 4: Mean prediction errors calculated using Eq. (9) of the GNN for each nucleus in the TENDL dataset. For better visualization, all plots are normalized between $\epsilon_{\rm min}$ and $0.03\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error over all plots. The three panels on the left side are based on VAE encoding while the three on the right on INR encoding. ARMAConv, GatedConv and GraphSAGE refer to the three types of GNNs considered in this work. The red square corresponds to the windowed test data: the GNN was not trained on it.
To quantify reconstruction performance, we use the Peak Signal-to-Noise Ratio (PSNR). The PSNR is defined as follows
@@
where $\text{MSE}=\frac{1}{n}\sum^{n}_{i=1}\|\mathbf{y}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2}$, $\hat{\mathbf{x}}_{i}$ represents the reconstructed input generated by either the VAE or the INR, while $\mathbf{y}_{i}$ denotes the original cross section inputs. Here, $n$ is the total number of testing samples. MAX represents the maximum possible value of the signal, which is equal to $1$. We evaluate each model using this metric on the entire test set. The VAE achieves a score of $46.68$, while the INR achieves a higher score of $51.39$. Overall, the performance of the INR is significantly better as this model is better able to precisely capture high-frequency information such as thresholds, for example the fact that the cross section for the $(n,2n)$ cross section is identically 0 unless the energy of the incident neutron exceeds the neutron separation energy.